d²y/dx² +2 dy/dx+y=e^-t lnt
1
Expert's answer
2020-12-02T19:01:51-0500
"solution"

Consider the differential equation


"y''+2y'+y=e^{-t}ln\ t"


The characteristic equation is


"m^2+2m+1=0\\
(m+1)^2=0\\
m+1=0\\
m=-1,-1"

So, the complimentary solution of the differential equation is


"y_c(t)=(c_1+c_2t)e^{-t}\\
y_c(t)=c_1e^{-t}+c_2te^{-t}\\"

Let "y_1(t)=e^{-t}, y_2(t)=te^{-t}"


The Wronskian of these two functions is

"w(y_1(t), y_2(t))=\begin{pmatrix}
e^{-t}& te^{-t} \\
-e^{-t} & e^{-t}-te^{-t} \\
\end{pmatrix}\\
\implies (e^{-2t}-te^{-2t})-(-te^{-2t})\\
\implies e^{-2t}"

And compare the given equation with "y''+p(t)y^t+q(t)y=g(t)" then "p(t)=2,\ q(t)=1, \ g(t)=e^t\ ln\ t"


Now the particular solution is "yp(t)=y_1u_1+y_2u_2" where

"u_1=-\intop\frac{y_2(t)g(t)}{w(y_1(t),y_2(t))}\delta t,\ u_2=\intop\frac{y_1(t)g(t)}{w(y_1(t),y_2(t))}\delta t,\\ 
u_1=-\intop\frac{(te^{-t})(e^{-t}ln\ t)}{e^{-2t}}\delta t,\ u_2=\intop\frac{(e^{-t})(e^{-t}ln\ t)}{e^{-2t}}\delta t,\\ 
u_1=-\intop(t\ ln\ t)\delta t,\ u_2=\intop(ln\ t)\delta t\\
u_1=-\frac{t^2}{4}(2\ ln\ t-1),\ u_2=t(ln\ t-1)"

Substitute "u_1, u_2" in the particular solution "yp(t)=y_1u_1+y_2u_2"

Then is

"yp(t)=(e^{-t})(-\frac{t^2}{4}(2\ ln\ t-1))+(te^{-t})(t(ln\ t-1))\\
\implies (-\frac{t^2}{4}(2\ ln\ t-1))+(ln\ t-1)t^2e^{-t})\\
\implies (-\frac{ln\ t}{2}+\frac14+ln\ t-1)t^2e^{-t}\\
\implies (\frac{ln\ t}{2}-\frac34)t^2e^{-t}\\"

Now, the general solution is

"y(t)=y_p(t)+y_c(t)"

"\therefore" The required solution is "y(t)=c_1e^{-t}+c_2te^{-t}+(\frac{ln\ t}{2}-\frac34)t^2e^{-t}"